OpenStudy (anonymous):

Simplify the following quotient of complex numbers into the form a + bi. -9+8i / 1+2i

OpenStudy (anonymous):

multiply top and bottom by conjugates (1-2i)

OpenStudy (anonymous):

ohhh i learned this this year (im 14)

myininaya (myininaya):

\[\frac{-9+8i}{1+2i} \cdot \frac{1-2i}{1-2i}\]

myininaya (myininaya):

now remember \[(a-b)(a+b)+a^2-b^2\]

OpenStudy (anonymous):

how is this figured?

OpenStudy (anonymous):

(-9 + 8i / 1 + 2i) * (1 -2i / 1 -2i)= -9 + 18i + 18i -16i^2 / 1^2 -4i^2 = -9 + 26i + 16 / 1 + 4 = 7 + 26i / 5

OpenStudy (anonymous):

idk what all of that means

OpenStudy (anonymous):

anyone want to check my work?

myininaya (myininaya):

\[\frac{-9+8i}{1+2i} \cdot \frac{1-2i}{1-2i}=\frac{-9+18i+8i-16i^2}{1-4i^2}=\frac{16-9+26i}{5}\] \[=\frac{7+26i}{5}\]

myininaya (myininaya):

gj